# POLAR: A Pessimistic Model-based Policy Learning Algorithm for Dynamic Treatment Regimes

### v1

@@ Abstract
Dynamic treatment regimes (DTRs) provide a principled framework for optimizing sequential decision-making in domains where decisions must adapt over time in response to individual trajectories, such as healthcare, education, and digital interventions. However, existing statistical methods often rely on strong positivity assumptions and lack robustness under partial data coverage, while offline reinforcement learning approaches typically focus on average training performance, lack statistical guarantees, and require solving complex optimization problems. To address these challenges, we propose POLAR, a novel pessimistic model-based policy learning algorithm for offline DTR optimization. POLAR estimates the transition dynamics from offline data and quantifies uncertainty for each history-action pair. A pessimistic penalty is then incorporated into the reward function to discourage actions with high uncertainty. Unlike many existing methods that focus on average training performance, POLAR directly targets the suboptimality of the final learned policy and offers theoretical guarantees, without relying on computationally intensive minimax or constrained optimization procedures. To the best of our knowledge, POLAR is the first model-based DTR method to provide both statistical and computational guarantees, including finite-sample bounds on policy suboptimality. Empirical results on both synthetic data and the MIMIC-III dataset demonstrate that POLAR outperforms state-of-the-art methods and yields near-optimal, history-aware treatment strategies.
@@ 1 Introduction
Despite these advances, both statistical DTR and offline RL approaches face limitations. Statistical DTR methods commonly rely on the positivity assumption, which requires that each action has a non-zero probability of being observed for any given patient history. However, this assumption is often violated in real-world observational data due to clinical guidelines and physician preferences, resulting in limited coverage of possible patient history-action combinations (Luckett et al.,, 2020). Offline RL methods, on the other hand, have been developed to address distribution shift in sequential decision-making but often lack statistical guarantees measured by sample size and the number of decision points. Additionally, offline RL algorithms typically rely on the Markov assumption, which is often violated in clinical settings where treatment decisions depend on the patient’s full history rather than just the current state, limiting the effectiveness of these methods. More importantly, many existing offline RL algorithms focus on the average policy performance during training or $\max_{0\leq t\leq T}$ metrics, rather than directly evaluating the suboptimality of the final learned policy (Chang et al.,, 2021; Uehara and Sun,, 2022). This distinction is critical in healthcare, where the final deployed policy directly impacts patient outcomes, making final policy suboptimality a more rigorous and clinically relevant evaluation criterion. Finally, many pessimistic offline RL methods require solving complex optimization problems, such as minimax or constrained formulations, which may limit their scalability and practical implementation in clinical decision-making contexts (Uehara and Sun,, 2022; Rashidinejad et al.,, 2022; Rigter et al.,, 2022).
To address these limitations, we develop POLAR, a novel pessimistic model-based policy learning algorithm for optimizing DTRs in offline settings. Unlike existing statistical and RL methods, POLAR explicitly accounts for uncertainty in poorly represented regions of the data, effectively addressing challenges related to partial data coverage and violations of the Markov assumption. POLAR first estimates transition dynamics from offline data and quantifies uncertainty for all history-action pairs. By incorporating a pessimistic penalty into the estimated reward function, POLAR discourages selecting actions associated with high uncertainty. Using this modified reward function and the estimated transition dynamics, POLAR iteratively updates the policy within an actor-critic framework, thus enabling efficient policy learning without solving complex optimization subproblems.
@@ 3.1 Overview of POLAR
Directly optimizing a policy using the estimated model $\widehat{M}=(\widehat{P},\widehat{r})$ can be suboptimal due to inaccuracies in $\widehat{P}$, especially in poorly represented state-action regions of the offline dataset. This arises from the distributional shift between the offline data and the distributions induced by policies other than the behavior policy. To mitigate this issue, we quantify the estimation uncertainty of $\widehat{P}_{k}$ by $\Gamma_{k}(h_{k},a_{k})$, and incorporate it into a penalized reward function. Specifically, we define a modified reward function $\widetilde{r}=\{\widetilde{r}_{k}\}_{k=1}^{K}$ as: $\widetilde{r}_{k}(h_{k},a_{k})=\widehat{r}_{k}(h_{k},a_{k})-\widetilde{c}_{k}\Gamma_{k}(h_{k},a_{k})$, where $\widetilde{c}_{k}\geq 0$ is a tuning multiplier that controls the degree of pessimism. This penalized reward function discourages the policy from visiting state-action pairs with high uncertainty. The modified DTR model, denoted as $\widetilde{M}=\left(\widehat{P},\widetilde{r}\right)$, combines the estimated transition dynamics $\widehat{P}$ with the pessimistically adjusted reward functions $\widetilde{r}$. POLAR optimizes the policy by maximizing the value function under this modified model. We will discuss the detailed construction of $\{\widehat{P}_{k}\}_{k=1}^{K}$, $\{\widehat{r}_{k}\}_{k=1}^{K}$ and $\{\Gamma_{k}\}_{k=1}^{K}$ in Section 3.2.
@@
Here $f_{k}(\theta_{k},h_{k},a_{k})$ is an importance-weighting function defined on $\Theta_{k}\times\mathcal{H}_{k}\times\mathcal{A}_{k}$, capturing the relative importance of action $a_{k}$ at stage $k$. The soft-max policy $\pi_{k}(\cdot|h_{k},\theta_{k})$, commonly used in the literature, defines a probability distribution over $\mathcal{A}_{k}$ and is parameterized by $\theta_{k}$. For clarity, we use $\pi(\theta)=(\pi_{1}(\theta_{1}),...,\pi_{K}(\theta_{K}))$ with $\theta=(\theta_{1},...,\theta_{K})$. The POLAR algorithm, summarized in Algorithm 1, follows a model-based actor-critic framework to optimize the policy under the modified DTR model within the policy class $\Pi$.
@@
3 Construct the modified DTR model $\widetilde{M}=(\widehat{P},\widetilde{r})$, with $\widetilde{r}_{k}=\widehat{r}_{k}-\widetilde{c}_{k}\Gamma_{k}$
@@ 3.2 The modified DTR
it holds with probability at least $1-\delta$ (with respect to $\mathbb{P}_{\mathcal{D}}$). That is, $\mathbb{P}_{\mathcal{D}}(\mathcal{E}_{k})\geq 1-\delta$ for each $k=1,...,K$.
@@
where $\phi_{k}:\mathcal{H}_{k}\times\mathcal{A}_{k}\to\mathbb{R}^{d_{k}^{(s)}}$ is a known bounded feature mapping, $F_{k}$ is a known distribution on $\mathbb{R}^{d_{k+1}^{(s)}}$, and $d_{k}^{(s)}$ is the dimension of the $k$-th state space $\mathcal{S}_{k}$. $W_{k}\in\mathbb{R}^{d_{k+1}^{(s)}\times d_{k}^{(s)}}$ is a weight matrix. The uncertainty quantifiers $\Gamma_{k}$ take the following form:
@@
Example 2: (Gaussian Process) Suppose the ground truth transition model follows a Gaussian Process (GP) structure: $s_{k+1}=g_{k}(h_{k},a_{k})+\varepsilon_{k}$, $\varepsilon_{k}\sim N(0,\sigma^{2}I)$, where $\varepsilon_{k}$ is Gaussian noise. Define $x:=(h_{k},a_{k})$ when $k$ is fixed. The function $g:\mathcal{X}\to\mathbb{R}^{d_{k+1}^{(s)}}$ is the mean transition dynamics, with ${d_{k+1}^{(s)}}$ denoting the dimension of the next state. We assume $g$ follows a GP prior: $g\sim GP(0,h(\cdot,\cdot))$, with zero mean and covariance function $h:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, where $h$ is a known symmetric positive semidefinite kernel.
Given observations $\{(x_{i},y_{i})\}_{i=1}^{n}$, the posterior distribution of $g$ remains a GP with mean function $\widehat{g}(\cdot)$ and covariance function $\widehat{h}(\cdot,\cdot).$ The uncertainty quantifier $\Gamma$ takes the form:
@@ 4 Theory
Terms $(a)$ and $(c)$ together capture the model shift error, i.e., the discrepancy between the true and estimated transition models. Since the modified DTR model is defined as $\widetilde{M}=\left(\widehat{P},\widetilde{r}\right)$ and the true model as $M^{*}=\left(P^{*},r\right)$, bounding $(a)$ and $(c)$ reduces to quantifying the difference between the estimated transition dynamics $\widehat{P}$ and the ground truth dynamics $P^{*}$, leveraging the construction of $\widetilde{r}$ in Section 3.2. Term $(b)$ captures the policy optimization error under the modified model $\widetilde{M}$. We present our main theoretical result in Section 4.1 and establish theoretical guarantees for two specific transition model cases in Section 4.2. All the main proofs and required technical assumptions are deferred to Supplement C. Auxiliary lemmas with their proofs are provided in Supplement D.
@@ 4.1 Main theoretical results
We begin by characterizing the function approximation error that arises at each iteration. At stage $k$, the policy update step (line 6 of Algorithm 1) can be viewed as approximating the target function $f_{k}(\theta_{k}^{(t)},\cdot,\cdot)+\eta_{k}^{(t)}Q^{\pi^{(t)}}_{k,\widetilde{M}}(\cdot,\cdot)$ using a new function $f_{k}(\theta_{k}^{(t+1)},\cdot,\cdot)\in\mathcal{F}_{k}=\{f_{k}(\theta_{k},\cdot,\cdot):\theta_{k}\in\Theta_{k}\}$. We define the corresponding approximation error at iteration $t$ as:
@@
For any $1\leq k\leq K$, there exists a quantity $G_{k}$ such that for any $10\leq t\leq T-$ 1,
@@
Suppose $\widetilde{M}=(\widehat{P},\widetilde{r})$ is the modified DTR model constructed in Algorithm 2. Under regularity conditions and specific hyperparameters detailed in Supplement C.1.2, Algorithm 2 returns a policy $\pi^{(T)}$ satisfying that:
@@ 4.2.1 Linear transition model
$\mathrm{Subopt}(\pi^{(T)};M^{*})\leq\mathcal{O}\left(\frac{K^{3}\sqrt{\log n}}{\sqrt{n}}+\frac{K^{5}}{n}\right)+\mathcal{O}\left(\frac{K}{\sqrt{T}}\right)+K\sqrt{T}\cdot\mathcal{O}_{p}\left(\left(\frac{\log m_{k}}{m_{k}}\right)^{\alpha_{1}}\right),$
@@
The first term controls the estimation error under finite offline sample size $n$, and vanishes as $n\to\infty$. The second term represents the policy optimization error, which vanishes as $T\to\infty$. The last term accounts for function approximation error in estimating the Q-function, which is controlled by the Monte Carlo sample size $m_{k}$. Based on this bound, we derive the following corollary, which quantifies the required offline sample size to guarantee an $\epsilon$-optimal policy in the limit as $T\to\infty$.
@@
Under the conditions of Theorem 2, with probability at least $1-K\delta$, the suboptimality gap satisfies $\mathrm{Subopt}(\pi^{(T)};M^{*})\leq\epsilon$ as $T,m_{k}\to\infty$, provided that the offline dataset size satisfies: $n=\widetilde{\Theta}\left(\frac{K^{6}}{\epsilon^{2}}\right).$
@@
Li et al., (2024) study offline RL in finite-horizon, tabular MDPs, whereas our setting focuses on offline learning in DTRs with continuous state spaces. Their method constructs an empirical MDP from offline data and uses Bernstein-style penalty terms to address transition uncertainty, achieving a sample complexity of $n=\widetilde{\Theta}\left(\frac{K^{3}}{\epsilon^{2}}\right)$ for an $\epsilon$-optimal policy. In contrast, our sample complexity bound in Corollary 1 is larger by a factor of $K^{3}$, i.e., $n=\widetilde{\Theta}\left(\frac{K^{6}}{\epsilon^{2}}\right)$, which is expected due to the structural differences between tabular MDPs and DTRs. In DTRs, each decision depends on the entire history of prior states and actions, leading to compounded estimation errors across stages and necessitating more data for accurate learning.
@@ 5.1 Simulation setup
We evaluate the performance of POLAR by simulation studies, and investigate the impact of pessimism on optimal policy estimation. We consider a 3-stage DTR setting, where each state-action trajectory takes the form $(s_{1},a_{1},s_{2},a_{2},s_{3},a_{3},s_{4})$. Assume that the state space is a two-dimensional continuous space, denoted as $s_{k}=(s_{k}^{1},s_{k}^{2})$, $k=1,\dots,4$. The action space is binary, with $a_{k}\in\{0,1\}$ for each stage. The initial state $s_{1}$ is generated uniformly from the unit square $[0,1]^{2}$. At each stage $k$, we define the true transition dynamics $P_{k}^{*}(\cdot\mid h_{k},a_{k})$ using a linear transition model: $s_{k+1}=W_{k}^{a}\phi(s_{k})+\varepsilon_{k}$, where $\phi(s_{k})=(1,s_{k}^{1},s_{k}^{2})$, $W_{k}^{a}$ is a $2\times 3$ matrix for each $k$ and $a$, and $\varepsilon_{k}=(\varepsilon_{k}^{1},\varepsilon_{k}^{2})$ is the noise term. The values of $W_{k}^{a}$ and the specification of $\varepsilon_{k}$ are provided in Supplement E. We assume that the reward is generated only at the final stage ($r_{1}=r_{2}=0$), and it is defined as $r=3.8\left[\left(\cos\left(-s_{3}^{1}\pi\right)+2\cos\left(s_{3}^{2}\pi\right)+s_{4}^{1}+2s_{4}^{2}\right)\left(1+a_{3}\right)-1.37\right]$.
@@
Based on the ground truth transition models and reward function, we first approximate the optimal policy $\pi^{*}$ using Dynamic Programming (Bellman et al.,, 1957). The implementation details are provided in Supplement F.1. We then construct the behavior policy $\pi_{b}$, which takes the optimal action $a^{*}$ given by $\pi^{*}$ with probability $p$, and takes the non-optimal action $1-a^{*}$ with probability $1-p$. Offline data are generated using $\pi_{b}$, consisting of $n$ trajectories. We explore various behavior policies by varying the probability parameter $p\in\{0.95,0.75,0.55\}$ and the offline sample size $n\in\{50,200,1000,5000,20000\}$. For each combination of $(p,n)$, we conduct $N=100$ repeated simulations.
@@
For each simulated dataset, we apply the proposed POLAR to estimate the optimal policy, resulting in the learned policy $\widehat{\pi}$. The value function of $\widehat{\pi}$ is then evaluated under the simulated true transition $P^{*}$ and reward function $r$ using Monte Carlo sampling. As the hyperparameter $c$ controls the degree of uncertainty penalization, we implement POLAR using a range of values of $c\in\{0,5,10,50,100\}$, where $c=0$ represents no pessimism. In our simulation study, $T=20$ steps proves sufficient for convergence. For comparison, we apply three alternative benchmark methods to each simulated dataset: (1) Q-learning for DTRs (DTR-Q), which estimates the state-action value function recursively (Murphy,, 2005), with implementation details provided in Supplement F.2; (2) a regression-based method for estimating optimal DTRs using the R package DTRreg (Wallace et al.,, 2024), which fits a sequence of outcome regression models to identify optimal actions at each stage; and (3) deep reinforcement learning with Double Q-learning (DDQN) (Van Hasselt et al.,, 2016), implemented in Python using deep neural networks to approximate the Q-function.

### v2

@@ Abstract
Dynamic treatment regimes (DTRs) provide a principled framework for optimizing sequential decision-making in domains where decisions must adapt over time in response to individual trajectories, such as healthcare, education, and digital interventions. However, existing statistical methods often rely on strong positivity assumptions and lack robustness under partial data coverage, while offline reinforcement learning approaches typically focus on average training performance, lack statistical guarantees, and require solving complex optimization problems. To address these challenges, we propose POLAR, a novel pessimistic model-based policy learning algorithm for offline DTR optimization. POLAR estimates the transition dynamics from offline data and quantifies uncertainty for each history-action pair. A pessimistic penalty is then incorporated into the reward function to discourage actions with high uncertainty. Unlike many existing methods that focus on average training performance or provide guarantees only for an oracle policy, POLAR directly targets the suboptimality of the final learned policy and offers theoretical guarantees, without relying on computationally intensive minimax or constrained optimization procedures. To the best of our knowledge, POLAR is the first model-based DTR method to provide both statistical and computational guarantees, including finite-sample bounds on policy suboptimality. Empirical results on both synthetic data and the MIMIC-III dataset demonstrate that POLAR outperforms state-of-the-art methods and yields near-optimal, history-aware treatment strategies.
@@ 1 Introduction
Despite these advances, both statistical DTR and offline RL approaches face limitations. Statistical DTR methods commonly rely on the positivity assumption, which requires that each action has a non-zero probability of being observed for any given patient history. However, this assumption is often violated in real-world observational data due to clinical guidelines and physician preferences, resulting in limited coverage of possible patient history-action combinations (Luckett et al.,, 2020). Offline RL methods, on the other hand, have been developed to address distribution shift in sequential decision-making but often lack statistical guarantees measured by sample size and the number of decision points. Additionally, offline RL algorithms typically rely on the Markov assumption, which is often violated in clinical settings where treatment decisions depend on the patient’s full history rather than just the current state, limiting the effectiveness of these methods. More importantly, many existing offline RL algorithms focus on metrics such as average policy performance during training or $\max_{0\leq t\leq T}$ bounds (Hong et al.,, 2024; Zanette et al.,, 2021), or on the performance of an oracle policy within a pessimistic model (Chang et al.,, 2021; Uehara and Sun,, 2022), rather than directly evaluating the suboptimality of the final learned policy. To address these limitations, we develop POLAR, a novel pessimistic model-based policy learning algorithm for optimizing DTRs in offline settings. Unlike existing statistical and RL methods, POLAR explicitly accounts for uncertainty in poorly represented regions of the data, effectively addressing challenges related to partial data coverage and violations of the Markov assumption. POLAR first estimates transition dynamics from offline data and quantifies uncertainty for all history-action pairs. By incorporating a pessimistic penalty into the estimated reward function, POLAR discourages selecting actions associated with high uncertainty. Using this modified reward function and the estimated transition dynamics, POLAR iteratively updates the policy within an actor-critic framework, thus enabling efficient policy learning without solving complex optimization subproblems.
@@ 3.1 Overview of POLAR
[…]
@@
Here $f_{k}(\theta_{k},h_{k},a_{k})$ is an importance-weighting function defined on $\Theta_{k}\times\mathcal{H}_{k}\times\mathcal{A}_{k}$, capturing the relative importance of action $a_{k}$ at stage $k$. The soft-max policy $\pi_{k}(\cdot|h_{k},\theta_{k})$, commonly used in the literature, defines a probability distribution over $\mathcal{A}_{k}$ and is parameterized by $\theta_{k}$. For clarity, we use $\pi(\theta)=(\pi_{1}(\theta_{1}),...,\pi_{K}(\theta_{K}))$ with $\theta=(\theta_{1},...,\theta_{K})$. The POLAR algorithm, summarized in Algorithm 1, follows a model-based actor-critic (Sutton and Barto,, 2018) framework to optimize the policy under the modified DTR model within the policy class $\Pi$. The actor is the parameterized policy $\pi(\theta)$, and the critic estimates the action-value function under the current policy. At each iteration $t$, the critic first computes the stage-wise $Q$-functions $Q^{\pi^{(t)}}_{k,\widetilde{M}}(h_{k},a_{k})$ under the modified model $\widetilde{M}$. The actor then updates $\theta$ so that actions with larger estimated $Q$ values are selected with higher probability under the soft-max policy. Full details of our actor-critic framework are given in Section 3.3.
@@
[…]
@@ 3.2 The modified DTR
it holds with probability at least $1-\delta$ (with respect to $\mathbb{P}_{\mathcal{D}}$). That is, $\mathbb{P}_{\mathcal{D}}(\mathcal{E}_{k})\geq 1-\delta$ for each k=1,,Kk=1,...,K.
@@
where $\phi_{k}:\mathcal{H}_{k}\times\mathcal{A}_{k}\to\mathbb{R}^{\dim(\phi_{k})}$ is a known bounded feature mapping, $F_{k}$ is a known distribution on $\mathbb{R}^{d_{k+1}^{(s)}}$, and $d_{k+1}^{(s)}$ is the dimension of the $(k+1)$-th state space $\mathcal{S}_{k+1}$. $W_{k}\in\mathbb{R}^{d_{k+1}^{(s)}\times\dim(\phi_{k})}$ is a weight matrix. The uncertainty quantifiers $\Gamma_{k}$ take the following form:
@@
Example 2: (Gaussian Process) Suppose the ground truth transition model follows a Gaussian Process (GP) structure: $s_{k+1}=g_{k}(h_{k},a_{k})+\varepsilon_{k}$, where $\varepsilon_{k}\sim\mathcal{N}(0,\sigma^{2}I)$. Define $x:=(h_{k},a_{k})$ when $k$ is fixed. The function $g:\mathcal{X}\to\mathbb{R}^{d_{k+1}^{(s)}}$ is the mean transition dynamics, with $d_{k+1}^{(s)}$} denoting the dimension of the next state. We assume $g$ follows a GP prior: $g\sim GP(0,h(\cdot,\cdot))$, with zero mean and covariance function $h:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, where $h$ is a known symmetric positive semidefinite kernel.
[…]
@@ 4 Theory
[…]
@@ 4.1 Main theoretical results
[…]
@@
For any $1\leq k\leq K$, there exists a quantity $G_{k}$ such that for any $0\leq t\leq T-1$,
@@
[…]
@@ 4.2.1 Linear transition model
$\mathrm{Subopt}(\pi^{(T)};M^{*})\leq\mathcal{O}\!\left(\frac{K^{2}\sqrt{\log n}}{\sqrt{n}}+\frac{K^{7/2}}{n}\right)+\mathcal{O}\!\left(\frac{K}{\sqrt{T}}\right)+K\sqrt{T}\cdot\mathcal{O}_{p}\!\left(\left(\frac{\log m_{k}}{m_{k}}\right)^{\alpha_{1}}\right),$
@@
The first term controls the estimation error under finite offline sample size $n$, and vanishes as $n\to\infty$. The second term represents the policy optimization error, which vanishes as $T\to\infty$. The last term accounts for function approximation error in estimating the $Q$-function, which is controlled by the Monte Carlo sample size $m_{k}$. In this bound, we focus on the asymptotic dependence on the sample size $n$, horizon $K$, and iteration count $T$. Problem-dependent constants including the coverage coefficients $\{C_{k}\}_{k=1}^{K}$, the uncertainty quantifier norms $\|\Gamma_{k}\|_{\infty}$, and the reward bounds $\|\overline{r}_{k}\|_{\infty}$ are absorbed into the Big-$O$ notation, as is standard in finite-sample analyses where such quantities are treated as fixed. Next, we establish a minimax lower bound, showing the rate is near-optimal.
Theorem 3 (Minimax Lower Bound of the Linear Transition Model).
There exist a universal constant $C_{\text{lower}}>0$ and a class $\mathcal{M}$ of DTR models with linear transition kernels such that the following holds. Given $n$ i.i.d. offline trajectories of length $K$, $\mathcal{D}=\big\{(s_{1}^{(i)},a_{1}^{(i)},r_{1}^{(i)},\dots,s_{K}^{(i)},a_{K}^{(i)},r_{K}^{(i)})\big\}_{i=1}^{n},$ for the final policy $\widehat{\pi}$ of any learning algorithm $\mathcal{A}$,
$\inf_{\mathcal{A}}\sup_{M\in\mathcal{M}}\mathbb{P}_{M}\!\left(V_{M}^{\pi^{*}}-V_{M}^{\widehat{\pi}}\;\geq\;C_{\text{lower}}\frac{K^{2}}{\sqrt{n}}\right)\;\geq\;\frac{1}{4}.$
Theorem 3 establishes a minimax lower bound for the same class of linear-transition DTR models, proving that any offline algorithm based on $n$ trajectories must incur a worst-case value error of at least $\Omega(K^{2}/\sqrt{n})$. When the number of policy optimization iterations $T\to\infty$, the dependence of our upper bound on $K$ and $n$ matches this lower bound up to a $\sqrt{\log n}$ factor. Therefore, Algorithm 2 is near-minimax optimal for this model class. Based on the above theorems, we derive the following corollary, which quantifies the required offline sample size to guarantee an $\epsilon$-optimal policy in the limit as $T\to\infty$.
@@
Under the conditions of Theorem 2, with probability at least $1-K\delta$, the suboptimality gap satisfies $\mathrm{Subopt}(\pi^{(T)};M^{*})\leq\epsilon$ as $T,m_{k}\to\infty$, provided that the offline dataset size satisfies $n=\widetilde{\Theta}\left(\frac{K^{4}}{\epsilon^{2}}\right).$
@@
Li et al., (2024) and Xiong et al., (2023) analyze offline RL in finite-horizon MDPs with a single transition kernel, under tabular and linear function approximation settings, respectively. They establish a sample complexity of $n=\widetilde{\Theta}\!\left(K^{3}/\varepsilon^{2}\right)$ for achieving an $\varepsilon$-optimal policy. In contrast, our sample complexity bound in Corollary 1 is larger by a factor of $K$, i.e., $n=\widetilde{\Theta}\left(\frac{K^{4}}{\epsilon^{2}}\right)$ and is minimax-optimal for the considered DTR linear transition model class. The difference in horizon dependence arises from the structural differences between the models. Our DTR model involves $K$ distinct stage-specific linear transition kernels, whose dimensions grow linearly with the horizon $K$. Therefore, the accumulation of stage-wise model shift errors inherently yields a larger dependence on the horizon.
@@ 5.1 Simulation setup
We evaluate the performance of POLAR by simulation studies, and investigate the impact of pessimism on optimal policy estimation. We consider a 3-stage DTR setting, where each state-action trajectory takes the form $(s_{1},a_{1},s_{2},a_{2},s_{3},a_{3},s_{4})$. Assume that the state space is a two-dimensional continuous space, denoted as $s_{k}=(s_{k}^{1},s_{k}^{2})$, $k=1,\dots,4$. The action space is binary, with $a_{k}\in\{0,1\}$ for each stage. The initial state $s_{1}$ is generated uniformly from the unit square $0,1]^{2}$. At each stage $k$, we define the true transition dynamics $P_{k}^{*}(\cdot\mid h_{k},a_{k})$ using a linear transition model: $s_{k+1}=W_{k}^{a}\phi(s_{k})+\varepsilon_{k}$, where $\phi(s_{k})=(1,s_{k}^{1},s_{k}^{2})$, $W_{k}^{a}$ is a $2\times 3$ matrix for each $k$ and $a$, and $\varepsilon_{k}=(\varepsilon_{k}^{1},\varepsilon_{k}^{2})$ is the noise term. The values of $W_{k}^{a}$ and the specification of $\varepsilon_{k}$ are provided in Supplement E. We assume that the reward is generated only at the final stage ($r_{1}=r_{2}=0$), and it is defined as $r=3.8\left[\left(\cos\left(-s_{3}^{1}\pi\right)+2\cos\left(s_{3}^{2}\pi\right)+s_{4}^{1}+2s_{4}^{2}\right)\left(1+a_{3}\right)-1.37\right]$.
@@
Based on the ground truth transition models and reward function, we first approximate the optimal policy $\pi^{*}$ using Dynamic Programming (Bellman et al.,, 1957). The implementation details are provided in Supplement F.1. We then construct the behavior policy $\pi_{b}$, which takes the optimal action $a^{*}$ given by $\pi^{*}$ with probability $p$, and takes the non-optimal action $1-a^{*}$ with probability 1p1-p. Offline data are generated using $\pi_{b}$, consisting of $n$ trajectories. We explore various behavior policies by varying the probability parameter $p\in\{0.95,0.75,0.55\}$ and the offline sample size $n\in\{50,200,1000,5000,20000\}$. For each combination of $(p,n)$, we conduct $N=100$ repeated simulations.
@@
For each simulated dataset, we apply the proposed POLAR to estimate the optimal policy, resulting in the learned policy $\widehat{\pi}$. The value function of $\widehat{\pi}$ is then evaluated under the simulated true transition $P^{*}$ and reward function $r$ using Monte Carlo sampling. As the hyperparameter $c$ controls the degree of uncertainty penalization, we implement POLAR using a range of values of $c\in\{0,5,10,50,100\}$, where $c=0$ represents no pessimism. In our simulation study, $T=20$ steps proves sufficient for convergence. For comparison, we apply five alternative methods to each simulated dataset: (1) Q-learning for DTRs (DTR-$Q$), which estimates the state-action value function recursively (Murphy,, 2005), with implementation details provided in Supplement F.2; (2) a regression-based method for estimating optimal DTRs using the R package DTRreg (Wallace et al.,, 2024), which fits a sequence of outcome regression models to identify optimal actions at each stage; (3) deep reinforcement learning with Double Q-learning (DDQN) (Van Hasselt et al.,, 2016), implemented in Python using deep neural networks to approximate the Q-function; (4) model-based offline policy optimization (MOPO) (Yu et al.,, 2020), which performs policy optimization using a learned dynamics model with an uncertainty-based penalty to discourage exploitation of out-of-distribution regions; and (5) model-based imitation learning from offline data (MILO) (Chang et al.,, 2021), which frames offline policy learning as a distribution-regularized imitation problem, constraining the policy to the behavior policy.